# Study of charged Lepton Flavor Violation in electron muon interactions

### v3

@@ 2 Physics processes and Monte Carlo simulation
In this study, we consider that there are not $Z-Z^{\prime}$ mixing, and the extra $Z^{\prime}$ current has the same gauge coupling and chiral strength as the standard model $Z^{0}$, but allows for the lepton flavor violation, similarly as Ref. ATLAS:2018mrn ; Li:2023lin . The coupling strength of the $Z^{\prime}$ and leptons can be described by a matrix $\lambda$ as Eq. 1
@@ 2.3 Event generation and simulation
There are some differences between the $\tau$ simulation and $e/\mu$. For those final states with $\tau$ in Tab.2, although they can go through any decay chains in reality, in this study we only consider the hadronic decay channels (about 60% of the total decay) and reconstruct it by the JET information.

### v4

@@ 2 Physics processes and Monte Carlo simulation
The full neutral current can be written as
$-\mathcal{L}_{\text{NC}}=eJ^{\mu}_{\text{em}}A_{\mu}+g_{1}J^{(1)\mu}Z^{0}_{(1),\mu}+g_{2}J^{(2)\mu}Z^{0}_{(2),\mu},$ (1)
where $Z^{0}_{(1)}$ and $Z^{0}_{(2)}$ respectively correspond to SM $SU(2)\times U(1)$ and extra $U(1)$, and both have the same coupling structure. In general, $Z^{0}_{(1)}$ and $Z^{0}_{(2)}$ will mix induced by the electroweak breaking, then related to the mass eigenstates $Z$ and $Z^{\prime}$ through an orthogonal transformation, while it is generally believed that the magnitude of this mixing angle is quite small, on the order of less than several $10^{-3}$ Langacker:2000ju .
In this study, we consider that there are not $Z-Z^{\prime}$ mixing, and the extra $Z^{\prime}$ current has the same gauge coupling and chiral strength as the SM $Z$, but allows for coupling with leptons of different generations, similarly as Ref. ATLAS:2018mrn ; Li:2023lin . The coupling strength of the $Z^{\prime}$ and leptons can be described by a matrix $\lambda$ as Eq. 2
@@ 2.3 Event generation and simulation
There are some differences between the $\tau$ simulation and $e/\mu$. For those final states with $\tau$ in Tab. 2, although they can go through any decay chains in reality, in this study we only consider the hadronic decay channels (about 60% of the total decay). The $\tau$ tagging efficiency is 80% with $p_{T}>10$ GeV, as defined in Delphes card.
@@ 3.1.1 Background study
Other remaining backgrounds, such as those arising from $\mu-\pi$ mis-identification, are not included in this study, since it is difficult to produce a significant contribution to the background level when taking into account both the mis-identification probability and the kinematic differences from signal.